OpenStudy (anonymous):

express (1-3j)/(3+4j) in the form a+jb where j is equal to the square root of -1

OpenStudy (anonymous):

$$-\frac{9}{25}-\frac{13}{25}i$$

OpenStudy (anonymous):

how does one do this though?

OpenStudy (anonymous):

If the denominator is of the form (a+bi), then multiply the top and bottom by (a-bi)

OpenStudy (anonymous):

Basically $$\frac{c+di}{a+bi}=\frac{(ac+bd)}{a^2+b^2}+\frac{(ad-bc)i}{a^2+b^2}$$

OpenStudy (anonymous):

For any c,d,a,b Just substitute your sht into there

OpenStudy (anonymous):

amazing thank you infinitely lad

OpenStudy (anonymous):

Here if you just want to get it done, type your equation into http://www.wolframalpha.com/

OpenStudy (anonymous):

Scroll down a couple lines